OpenStudy (dls):

What is the square root of 36?

OpenStudy (anonymous):

6

OpenStudy (dls):

\[6 or \pm6\]

OpenStudy (unklerhaukus):

\[\sqrt{36}= ±6\]

OpenStudy (dls):

How can square root be negative sire?

OpenStudy (unklerhaukus):

\[-6\times-6=36\]

OpenStudy (dls):

no

OpenStudy (unklerhaukus):

what is \[-6\times-6=\]

OpenStudy (dls):

square root is always positive and what you wrote is a different thing

OpenStudy (unklerhaukus):

what have i done?

OpenStudy (dls):

x=36 has only ONE DEGREE so how can it have two values sire?

OpenStudy (unklerhaukus):

maybe i am confused

OpenStudy (helder_edwin):

\[ \large \sqrt{36}=6 \]

jimthompson5910 (jim_thompson5910):

The square root function is just that...a function

jimthompson5910 (jim_thompson5910):

one number in -----> one number out

OpenStudy (anonymous):

Square root is not always positive. One can agree to only think about the positive square root, but that is just a convenience.

OpenStudy (dls):

but one degree can have only one possible solution

OpenStudy (dls):

If it was square then it would've been 2 degree then both plus and minus signs would've been there

OpenStudy (anonymous):

6 can u help me with mine pleeaazzzze

OpenStudy (radar):

As was mentioned above every positive number has two square roots (and so does wikipedia. The positive square root is denoted as the principal square root but the negative square root is also valid as (-6)(-6)=+36 or\[(-6)^{2}=36\]

OpenStudy (radar):

One number in (6) one number out 36 one number in (-6) one number out 36

jimthompson5910 (jim_thompson5910):

The square root of 36 is 6 (and only 6) The solutions to x^2 = 36 are x = 6 or x = -6 I think that's where the confusion lies

OpenStudy (anonymous):

i was always taught it was +6 or -6 that is why the in the quadratic formula you have to solve both ways